OpenStudy (anonymous):

Sam’s plasma TV has a 14 in by 18 in screen with a plastic trim border. The total area of the TV and border is 480 in2. How wide is the trim around the TV?

OpenStudy (anonymous):

Area(TV and Border)=480 in2 480=(14+x)(18+x) where x is the trim width 480=252+14x+18x+x^2 0=-228+32x+x^2 0=(x-6)(x+38) x=6 and x=-38 Since the width can't be negative it is x=6